# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ 1.3 Layout of the paper
In Section 2, we introduce the notion of the reshaping parameters with a motivating example. In Section 3, we describe the data-driven problem and the algorithm for learning reshaping parameters, and prove its convergence. In Section 4, we discuss the probabilistic guarantees implied by the aforementioned formulations. In Section 5, we present various numerical examples problems. For completeness, we give example convex reformulations of robust optimization problems for common uncertainty sets, with our reshaped parameters, in the appendices.
@@ 2 The robust optimization problem
We consider a family of parametric robust optimization (RO) problems, indexed by the family parameter $y\in{\mbox{\bf R}}^{p}$, where each instance takes the form:
@@
$x(\theta,y)\in\begin{array}[t]{ll}\mathop{\rm argmin}&f(x)\\ \mbox{subject to}&g(x,u,y)\leq 0\quad\forall u\in{\mathcal{U}}(\theta),\end{array}$ (1)
@@
where $x\in{\mbox{\bf R}}^{n}$ is the optimization variable, $u\in{\mbox{\bf R}}^{m}$ is the uncertain parameter, $f:{\mbox{\bf R}}^{n}\to{\mbox{\bf R}}$ is the convex objective function, and $g:{\mbox{\bf R}}^{n}\times{\mbox{\bf R}}^{m}\times{\mbox{\bf R}}^{p}\to{\mbox{\bf R}}$ is the uncertain constraint which we assume to be the maximum of functions concave in $u$ and convex in $x$ and $y$e.g., $g(x,u,y)=\max_{l=1,\dots,L}g_{l}(x,u,y)$. This structure allows us to express $L$ joint uncertain constraints using the maximum of each component $l$. Problem (1) guarantees that, for a given $y$, the solution $x(\theta,y)$ satisfies the constraints for any value of the uncertain parameter $u$ in ${\mathcal{U}}(\theta)\subseteq{\mbox{\bf R}}^{m}$, a convex uncertainty set that depends on the set parameter $\theta\in{\mbox{\bf R}}^{q}$. For example, $\theta$ can represent the center, size, and shape of ${\mathcal{U}}(\theta)$. In this setting, the family parameter $y$ represents information available when we make our decisions (e.g., current allocations in portfolio optimization). The uncertain parameter $u$, however, is not known (e.g., future returns in portfolio optimization). Importantly, we do not assume that $y$ affects $u$, which is common in several practical applications (e.g., when our portfolio allocations do no affect future market returns).
Here, the set ${\mathcal{U}}(\theta)$ does not depend on individual $y$ as this would require either learning a different uncertainty set for every instance of the family, which may be costly, or learning a mapping from $y$ to the uncertainty set (as in conditional RO (Sadana et al.,, 2023)), which may be hard to interpret. Instead, we construct a fixed set that works for all realizations of $y$ encountered in target application.
@@
Our goal is to find a set parameter $\theta$ in a decision-focused manner, i.e., such that the solutions $x(\theta,y)$ perform well in terms of objective value and constraint satisfaction guarantees across the family of problems. Specifically, we choose the uncertainty set ${\mathcal{U}}(\theta)$ such that the solution $x(\theta,y)$ from (1) implies that the uncertain constraint is satisfied with probability $1-\eta$ with $\eta>0$, across all instances of problem (1):
${\mathbf{P}}_{(u,y)}(g(x(\theta,y),u,y)\leq 0)\geq 1-\eta.$ (2)
@@
Here, ${\mathbf{P}}_{y}$ is the distribution of $y$ with domain $\mathcal{D}_{y}\subseteq{\mbox{\bf R}}^{p}$, ${\mathbf{P}}_{u}$ is the distribution of $u$ with domain $\mathcal{D}_{u}\subseteq{\mbox{\bf R}}^{m}$, and ${\mathbf{P}}_{(u,y)}$ is their joint distribution, all of which are unknown. We note that this is an aggregate probabilistic guarantee for the parametric family of problems, and not conditional guarantee which holds for all instances $y$. Whereas such conditional guarantee (Chenreddy and Delage,, 2024) are perhaps more desirable they require stringent assumptions on the distribution ${\mathbf{P}}_{y}$ which are typically hard to verify in practice. We can formulate constraint (2) in terms of the corresponding value at risk, i.e.,
$\mathop{\bf VaR}(g(x(\theta,y),u,y),\eta)=\inf\{\gamma\mid{\mathbf{P}}_{(u,y)}(g(x(\theta,y),u,y)\leq\gamma)\geq 1-\eta\}\leq 0.$
Unfortunately, the value at risk function often leads to intractable formulations (Uryasev and Rockafellar,, 2001). To get a tractable approximation, we can adopt the conditional value at risk (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002), defined as
$\mathop{\bf CVaR}(g(x(\theta,y),u,y),\eta)=\inf_{\alpha}\left\{\frac{{\mbox{\bf E}}_{(u,y)}(g(x(\theta,y),u,y)-\alpha)_{+}}{\eta}+\alpha\right\},$ (3)
where $(a)_{+}=\max\{a,0\}$. It is well known (Uryasev and Rockafellar,, 2001) that, for any $x\in{\mbox{\bf R}}^{n}$, the relationship between these probabilistic guarantees of constraint satisfaction is
$\mathop{\bf CVaR}(g(x,u,y),\eta)\leq 0~{}~{}\Longrightarrow~{}~{}\mathop{\bf VaR}(g(x,u,y),\eta)\leq 0~{}~{}\Longleftrightarrow~{}~{}{\mathbf{P}}_{(u,y)}(g(x,u,y)\leq 0)\geq 1-\eta.$ (4)
@@
Therefore, if the solution $x(\theta,y)$ of (1) satisfies $\mathop{\bf CVaR}(g(x(\theta,y),u,y),\eta)\leq 0$, we have the desired probabilistic guarantee (2). This constraint, together with the goal of minimizing the expected objective value, are constructed into an evaluation metric for updating $\theta$; see Figure 1. Typically, by constructing uncertainty sets that contain at least $1-\eta$ probability mass, i.e.${\mathbf{P}}_{u}(u\in{\mathcal{U}}(\theta))\geq 1-\eta$, we can ensure that any feasible solution of (1) implies a probabilistic guarantee (2). However, this sufficient condition may lead to overly conservative solutions (Ben-Tal et al.,, 2009, page 34), as it works for any RO problem and any constraint using the uncertainty set ${\mathcal{U}}(\theta)$. In this work, instead, we avoid such conservatism by taking into account the cost of the robust solutions $x(\theta,y)$ while constructing the uncertainty sets.
@@
Figure 1: Our high-level procedure. We evaluate the performance of the decision $x(\theta,y)$ in terms of the expected objective value and constraint satisfaction. The evalution metric is defined in Section 3.2.
@@ 2.1 Motivating example
We consider a 2-product newsvendor problem where each day, the vendor orders $x\in{\mbox{\bf R}}^{2}$ products at price $k\in{\mbox{\bf R}}^{2}$, and we allow fractional values for simplicity. The products are sold at price $p\in{\mbox{\bf R}}^{2}$, until either the demand $u\in{\mbox{\bf R}}^{2}$ or inventory $x$ is exhausted. The objective function to minimize is the sum of the ordering cost minus the revenue, $f(x,u,k,p)=k^{T}x-p^{T}\min\{x,u\}=k^{T}x+\max(-p_{1}x_{1}-p_{2}x_{2},-p_{1}x_{1}-p_{2}u_{2},-p_{1}u_{1}-p_{2}x_{2},-p_{1}u_{1}-p_{2}u_{2}).$ We obtain the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t\\ \mbox{subject to}&f(x,u,k,p)\leq t\quad\forall u\in{\mathcal{U}}(\theta)\\ &x\geq 0.\end{array}$ (5)
We parametrize the problem (5) by $y=(k,p)$, the buying and selling prices of the products, and would like to construct uncertainty sets without knowing the distributions of $y$ and $u$.
@@
The uncertain parameter $u$ is distributed as a log-normal distribution, with parameters
$\mu=\left[\begin{array}[]{l}1.1\\ 1.7\end{array}\right],\quad\Sigma=\left[\begin{array}[]{ll}\phantom{-}0.6&-0.3\\ -0.3&\phantom{-}0.1\end{array}\right],$
and we have access to $50$ independent samples. We generate 10 values of $(k,p)$, where each $k=[4,5]+\delta_{1}$ and $p=k+\max\{0,\delta_{2}\}$, with $\delta_{1},\delta_{2}$ standard normal random variables. We repeat the data generation for $u$ and $y$ detailed above, for 15 repetitions of the experiment. Below, we show the uncertainty sets obtained for one such repetition.
Figure 2: Top (bottom) row: mean-variance (reshaped) uncertainty set, depicted in red, corresponding to the empirical probability of constraint violation $\hat{\eta}$. White dots: training data points. Green dot: training data mean. Middle row: out-of-sample trade-off curves for the objective value vs. empirical probability of constraint violation, averaged across all problem instances and all repetitions of the experiment. The shaded regions represent the 0.3 to 0.7 quantiles of the repetitions. Each marker corresponds to an $\rho$ value. The uncertainty sets depicted above and below achieves the probabilities of constraint violation $\hat{\eta}$ closest to the red dotted lines, for one repetition of the experiment.
Mean-Variance uncertainty set.
@@
Standard methods in RO methods construct uncertainty sets based on the empirical mean $\hat{\mu}$ and covariance $\hat{\Sigma}$ of the uncertainty,
${\mathcal{U}}{(\theta)}=\{u=\hat{\mu}+\hat{\Sigma}^{1/2}z\mid\|z\|_{2}\leq\rho\}=\{b^{\rm mv}+A^{\rm mv}z\mid\|z\|_{2}\leq\rho\},$ (6)
@@
where $\theta=(A^{\rm mv},b^{\rm mv},\rho)$. The parameter $\rho$ represents the size of the uncertainty set, and, for one set of data, we have
$A^{\rm mv}=\hat{\Sigma}^{1/2}=\left[\begin{array}[]{ll}\phantom{-}3.08&-0.83\\ -0.83&\phantom{-}1.49\end{array}\right],\quad b^{\rm mv}=\hat{\mu}=\left[\begin{array}[]{l}3.96\\ 5.82\end{array}\right].$
Decision-focused (reshaped) uncertainty set.
Consider now the same set as in (6) with different shape, i.e., ${\mathcal{U}}(\theta)=\{u=b^{\rm re}+A^{\rm re}z\mid\|z\|_{2}\leq\rho\}$, where $\theta=(A^{\rm re},b^{\rm re},\rho)$ and
$A^{\rm re}=\left[\begin{array}[]{ll}0.12&2.05\\ 2.09&0.62\end{array}\right],\quad b^{\rm re}=\left[\begin{array}[]{l}2.7\\ 4.8\end{array}\right].$
@@
In Figure 2, we compare the two uncertainty sets where we vary $\rho$ while $A^{\rm st},b^{\rm st}$ and $A^{\rm re},b^{\rm re}$ are fixed. For each value of $\rho$ we compare the average out-of-sample objective value at the optimizers $x(\theta,y)$ against the out-of-sample empirical probability of constraint violation. The average is taken across problem instances $y$’s, testing datasets of $N=450$ uncertainty realizations, and over all repetitions of the experiment. We note that the reshaped uncertainty set achieves data-driven solutions that give, on average, better trade-offs. While the mean-variance uncertainty set conforms to the shape of data, it might be conservative because it does not take into account the structure of the optimization problem. In contrast, the reshaped uncertainty set gives better (lower) out-of-sample costs for the same probability of constraint satisfaction, even though it does not conform as well to the shape of the data, and covers a lower percentage of the data points.
In the next sections, we introduce an automatic technique to find such sets with better performance, by tuning their shape and size such that the problem objective is minimized while guaranteeing constraint satisfaction across the family of parametric RO problems (1).

### v5

@@ 1.3 Layout of the paper
In Section 2, we introduce the notion of the decision-focused set with a motivating example. In Section 3, we describe the data-driven algorithm and prove its convergence, and in Section 4, provide probabilistic guarantees. In Section 5, we discuss hyperparameter tuning, and in Section 6, present various numerical examples problems. For completeness, we give example convex reformulations of robust optimization problems for common uncertainty sets, with our shaped parameters, in the appendices.
@@ 2 The robust optimization problem
We consider a class of parametric robust optimization (RO) problems, indexed by the context parameter $x\in{\mathbf{R}}^{p}$, where each instance takes the form:
@@
$z(\theta,x)\in\begin{array}[t]{ll}\mbox{argmin}&f(z)\\ \mbox{subject to}&g(z,u,x)\leq 0\quad\forall u\in{\mathcal{U}}(\theta,x),\end{array}$ (1)
where $z\in{\mathbf{R}}^{n}$ is the optimization variable, $u\in{\mathbf{R}}^{m}$ is the uncertain parameter, $f:{\mathbf{R}}^{n}\to{\mathbf{R}}$ is the convex objective function, and $g:{\mathbf{R}}^{n}\times{\mathbf{R}}^{m}\times{\mathbf{R}}^{p}\to{\mathbf{R}}$ is the uncertain constraint which we assume to be the maximum of functions concave in $u$ and convex in $z$ and $y$e.g., $g(z,u,x)=\max_{l=1,\dots,L}g_{l}(z,u,x)$. This structure allows us to express $L$ joint uncertain constraints using the maximum of each component $l$. We note that linear adjustable RO problems with affine decision rules also fall within this class.
@@
Problem (1) guarantees that, for a given $x$, the solution $z(\theta,x)$ satisfies the constraints for any value of the uncertain parameter $u$ in ${\mathcal{U}}(\theta,x)\subseteq{\mathbf{R}}^{m}$, a convex contextual uncertainty set that depends on the set parameter $\theta\in{\mathbf{R}}^{q}$. For example, $\theta$ can represent the mapping between $x$ and the center, size, and shape of ${\mathcal{U}}(\theta,x)$. In this setting, the context parameter $x$ represents information available when we make our decisions (e.g., current allocations in portfolio optimization, the price of the products to sell in inventory problems). The uncertain parameter $u$, however, is not known (e.g., future returns in portfolio optimization). The context parameter $x$ is allowed to affect the uncertain parameter $u$.
@@
Our goal is to find a set parameter $\theta$ in a decision-focused manner, i.e., such that the solutions $z(\theta,x)$ perform well in terms of a performance metric and satisfy the constraint satisfaction guarantees across the problem class. Specifically, we choose the uncertainty set ${\mathcal{U}}(\theta,x)$ such that the solutions $z(\theta,x)$ from (1) imply that the uncertain constraint is satisfied with probability $1-\eta$ with $\eta>0$, across all instances of problem (1):
${\mathbf{P}}_{(u,x)}(g(z(\theta,x),u,x)\leq 0)\geq 1-\eta.$ (2)
@@
Here, ${\mathbf{P}}_{x}$ is the distribution of $x$ with domain $\mathcal{D}_{x}\subseteq{\mathbf{R}}^{p}$, ${\mathbf{P}}_{u}$ is the distribution of $u$ with domain $\mathcal{D}_{u}\subseteq{\mathbf{R}}^{m}$, and ${\mathbf{P}}_{(u,x)}$ is their joint distribution, all of which are unknown. Importantly, $u$ and $x$ can be correlated. We note that this is an aggregate probabilistic guarantee for the parametric class of problems, and not a conditional guarantee which holds for all instances $x$. Whereas such a conditional guarantee is perhaps more desirable, it requires stringent assumptions on the distribution ${\mathbf{P}}_{x}$, which are typically hard to verify in practice (Vovk,, 2012) (Angelopoulos and Bates,, 2023, Section 3.1).
@@
It is well known that by constructing uncertainty sets that contain at least $1-\eta$ probability mass, i.e.${\mathbf{P}}_{(u,x)}(u\in{\mathcal{U}}(\theta,x))\geq 1-\eta$, we can ensure that any feasible solution of (1) implies a probabilistic guarantee (2). However, this sufficient condition may lead to overly conservative solutions (Ben-Tal et al.,, 2009, page 34), as it works for any RO problem and any constraint using the uncertainty set ${\mathcal{U}}(\theta,x)$. In this work, instead, we avoid such conservatism by taking into account the average cost ${\mbox{\bf E}}_{(u,x)}\left(\Phi(z(\theta,x),u,x)\right)$ of the robust solutions $z(\theta,x)$ while constructing the uncertainty sets, where $\Phi(\cdot)$ is a given performance metric. For problems with only constraint uncertainty, this performance metric can be taken as simply the objective function, $f(\cdot)$. For problems with objective uncertainty, we delay the discussion to Section 2.1 below. Overall, we aim to find the best uncertainty set parametrization $\theta$ that satisfies the probabilistic constraint and also performs well on average. While uncertainty sets constructed from constraint information has been explored in recent literature (Bertsimas and Sim,, 2004; Bertsimas et al.,, 2018), systematic customization of a large number of uncertainty set parameters in a contextual and decision-focused manner is a novel topic. With these goals in mind, the ideal problem we would like to solve is
[…]
where $z(\theta,x)$ is the solution of the RO problem (1) for each $x$. Unfortunately, as the distributions are unknown and the chance constraint is intractable, this problem cannot be solved directly. A common reformulation technique is to express the chance constraint (2) in terms of the corresponding value at risk, i.e.,
$\mathop{\bf VaR}(g(z(\theta,x),u,x),\eta)=\inf\{\gamma\mid{\mathbf{P}}_{(u,x)}(g(z(\theta,x),u,x)\leq\gamma)\geq 1-\eta\}\leq 0.$
However, the value at risk function often also leads to intractable formulations (Uryasev and Rockafellar,, 2001). To get a tractable approximation, we can adopt the conditional value at risk (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002), defined as
$\mathop{\bf CVaR}(g(z(\theta,x),u,x),\eta)=\inf_{\alpha}\left\{\frac{{\mbox{\bf E}}_{(u,x)}(g(z(\theta,x),u,x)-\alpha)_{+}}{\eta}+\alpha\right\},$ (4)
where $(a)_{+}=\max\{a,0\}$. It is well known (Uryasev and Rockafellar,, 2001) that, for any $z\in{\mathbf{R}}^{n}$, the relationship between these probabilistic guarantees of constraint satisfaction is
$\mathop{\bf CVaR}(g(z,u,x),\eta)\leq 0~{}~{}\Longrightarrow~{}~{}\mathop{\bf VaR}(g(z,u,x),\eta)\leq 0~{}~{}\Longleftrightarrow~{}~{}{\mathbf{P}}_{(u,x)}(g(z,u,x)\leq 0)\geq 1-\eta.$ (5)
Therefore, if the solution $z(\theta,x)$ of (1) satisfies $\mathop{\bf CVaR}(g(z(\theta,x),u,x),\eta)\leq 0$, we have the desired probabilistic guarantee (2). In fact, as the $\mathop{\bf CVaR}$ constraint is stronger than the $\bf{VaR}$ constraint, ensuring its satisfaction might result in more desirable solutions, especially when, on top of the threshold worst-case value, the decision-maker is concerned with the magnitude of all the worst-case values beyond the threshold (Sarykalin et al.,, 2014). Furthermore, this formulation should not favor neglecting certain contexts $x$ (such that they suffer constraint violations of a significant magnitude) in exchange for better performance for other contexts, since that would result in a high $\mathop{\bf CVaR}$.
To solve the bi-level problem (3), we thus employ the $\mathop{\bf CVaR}$ approximation for the chance constraint, and use a stochastic gradient-based approach to learn an optimal parametrization $\theta$ of the uncertainty set. For the high-level architecture, see Figure 1.
@@
Figure 1: Our high-level procedure. We evaluate the performance of the decision $z(\theta,x)$ in terms of the expected objective value and constraint satisfaction. The evaluation metric is defined in Section 3.2.
2.1 Objective uncertainty and average performance
In the robust formulation (1), we assumed only uncertain constraints. However, robust objective functions can be easily rewritten in the given form, by introducing an epigraph variable, i.e., $t$, along with the uncertain constraint
$g_{0}(z,t,u,x)=f(z,u,x)-t\leq 0.$
The $\mathop{\bf CVaR}$ constraint $\mathop{\bf CVaR}(g_{0}(z(\theta,x),t(\theta,x),u,x),\eta)\leq 0$ then implies the probabilistic guarantee
${\mathbf{P}}_{(u,x)}(f(z(\theta,x),u,x)\leq t(\theta,x))\geq 1-\eta,$
such that $t(\theta,x)$ is an upper bound on the $\eta$-quantile worst-case objective. By minimizing ${\mbox{\bf E}}_{x}\left(t(\theta,x)\right)$ in the outer level of the bi-level problem (3), we optimize the $\eta$-quantile worst-case performance, averaged over the contexts. In many applications, however, in addition to the worst-case performance, we also wish to optimize the average-case performance. Related is the recent development on Pareto-optimal solutions (Iancu and Trichakis,, 2014), where the authors identify robust solutions that perform well not only in the worst-case. In this work, by defining a weighted performance metric, we can also solve for uncertainty sets that perform well in the average case. We consolidate the two objectives (minimizing worst-case and average-case performance) through the following formulation
${\mbox{\bf E}}_{(u,x)}\left(\gamma t(\theta,x)+(1-\gamma)f(z(\theta,x),u,x)\right),$ (6)
where $\gamma$ is a tunable hyper-parameter. In terms of (3), the performance metric is thus defined as $\Phi(z(\theta,x),u,x)=\gamma t(\theta,x)+(1-\gamma)f(z(\theta,x),u,x).$ Controlling the value of $\gamma$ allows us to trade-off the two objectives, and identify a solution that provides numerically the best of both worlds.
@@ 2.2 Motivating example
We demonstrate through a motivating example the average-case gains we could obtain through a decision-focused uncertainty set, even when the worst-case objective value cannot be further improved. We consider a 2-product newsvendor problem where each day, the vendor orders $z\in{\mathbf{R}}^{2}$ products at price $k\in{\mathbf{R}}^{2}$, and we allow fractional values for simplicity. The products are sold at price $p\in{\mathbf{R}}^{2}$, until either the demand $u\in{\mathbf{R}}^{2}$ or inventory $z$ is exhausted. The objective function to minimize is the sum of the ordering cost minus the revenue, $f(z,u,k,p)=k^{T}z-p^{T}\min\{z,u\}=k^{T}z+\max(-p_{1}z_{1}-p_{2}z_{2},-p_{1}z_{1}-p_{2}u_{2},-p_{1}u_{1}-p_{2}z_{2},-p_{1}u_{1}-p_{2}u_{2}).$ Using the epigraph form as described in Section 2.1, we obtain the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t\\ \mbox{subject to}&f(z,u,k,p)\leq t\quad\forall u\in{\mathcal{U}}(\theta,k,p)\\ &z\geq 0.\end{array}$ (7)
We parametrize the problem (7) with context parameters $x=(k,p)$, the buying and selling prices of the products, and would like to construct uncertainty sets without knowing the distributions of $x$ and $u$.
@@
We consider 20 unique sets of contexts, each repeated 100 times, to generate 2000 total data-samples. Each unique $k=(4,5)+\delta_{1}$ and $p=k+\max\{0,\delta_{2}\}$, with $\delta_{1},\delta_{2}$ normal random variables with mean 0 and variance 3. The uncertain parameter $u$ is distributed as a contextual normal distribution, with parameters $\mu=(3,4)-0.1p-0.2k,~{}\Sigma=I$, where the mean depends on the context. In other words, the demands of the products are inversely related to their prices. Overall, we generate 2000 uncertainty-context pairs $(u_{i},x_{i})$. We split the 2000 samples into 600 training data-points, 400 validation data-points, and 1000 testing data-points.
Methods.
As this is a contextual problem with joint uncertain constraints, it cannot be solved by methods that consider only concave objective uncertainty. Furthermore, for custom uncertain-constraint-based uncertainty sets, solving (7) would require a cutting plane algorithm for each uncertain constraint (each component of the maximum), and a separate procedure for each context. For this example, we thus compare our method against a mean-variance uncertainty set, and a contextual mean-variance uncertainty set, where such procedures are not required. We delay a more comprehensive comparison against other baselines to the numerical examples in Section 6. For each method, we tune the uncertainty set for at most a 10% probability of constraint violation on the validation dataset, and record the corresponding test-set values for the probability of constraint violation, 10% worst-case objective, and average objective. We repeat the data generation for $u$ and the uncertainty set training and selection for 10 runs of the experiment, and report the averaged values. The probability of constraint violation is measured as $\hat{{\mathbf{P}}}_{(u,x)}(f(z(\theta,x),u,x)\leq t(\theta,x)),$ where $\hat{{\mathbf{P}}}_{(u,x)}$ is the empirical distribution over the considered dataset. The worst-case and average-case objectives are likewise computed using the uncertain objective function $f(\cdot)$ over the considered datasets.
@@ Methods.
  • Mean-Variance (MV) uncertainty set. Standard methods in RO methods construct uncertainty sets based on the empirical mean $\hat{\mu}$ and covariance $\hat{\Sigma}$ of the uncertainty,
    ${\mathcal{U}}{(\theta)}=\{u=\hat{\mu}+\hat{\Sigma}^{1/2}v\mid\|v\|_{2}\leq\rho\}=\{b^{\rm mv}+A^{\rm mv}v\mid\|v\|_{2}\leq\rho\},$ (8)
@@
    where $\theta=(A^{\rm mv},b^{\rm mv},\rho)$. We tune only $\rho$, which represents the size of the uncertainty set.
  • Contextual Mean-Variance (CMV) uncertainty set. Consider now a contextual version of (8), i.e., ${\mathcal{U}}(\theta,x)=\{u=b^{\rm cmv}(x)+A^{\rm cmv}(x)v\mid\|v\|_{2}\leq\rho\}$, where $b^{\rm cmv}(x)$ is a linear mapping of the context $x$ to the contextual mean of the uncertainty, and $A^{\rm cmv}(x)$ is based on the covariance matrix of the $k$ uncertainty realizations associated with the $k$ nearest neighbors of $x$i.e.,
    $A^{\rm cmv}=\hat{\Sigma}^{1/2}(x),\quad b^{\rm cmv}(x)=W_{b}x+h_{b}.$ (9)
    In a predict-then-optimize manner, without considering the downstream optimization problem, we use the training samples of $(u_{i},x_{i})$ to train the weights and bias $W_{b}\in{\mathbf{R}}^{m,p},h_{b}\in{\mathbf{R}}^{m}$ of the mapping, and establish a $k$-nearest neighbor fit for each context; see Messoudi et al., (2022) for further details. We use 10% of the training data for the nearest neighbors. After the pre-training, the parameters of the uncertainty set become $\theta=(W_{b},h_{b},\hat{\Sigma}^{1/2}(\cdot),\rho)$, where only $\rho$ is subject to adjustments based on the validation dataset.
  • Decision-focused (LRO) uncertainty set. Consider now a reshaped ellipsoidal set
    ${\mathcal{U}}(\theta)=\{u=b^{\rm lro}(x)+A^{\rm lro}(x)v\mid\|v\|_{2}\leq\rho\},$
    where both $A^{\rm lro}(x)$ and $b^{\rm lro}(x)$ are linear mappings of the context $x$, i.e.,
    $A^{\rm lro}(x)=W_{A}x+h_{A},\quad b^{\rm lro}(x)=W_{b}x+h_{b},$ (10)
    where $W_{A}\in{\mathbf{R}}^{m,m,p},~{}h_{A}\in{\mathbf{R}}^{m,p},~{}W_{b}\in{\mathbf{R}}^{m,p},~{}h_{b}\in{\mathbf{R}}^{m}$. Note that the mappings for $A^{\rm lro}$ are tensors. The parameters of the uncertainty set are $\theta=(W_{A},W_{b},h_{A},h_{b},\rho)$, the first four of which are selected using an end-to-end framework, and the radius is tuned using the validation dataset. The performance metric we minimize is of the form (6).
Table 1: Out-of-sample values for the three methods, with a range between the 0.25 and 0.75 quantiles.
LRO MV CMV
$\hat{\eta}$ $0.0995$ $0.094$ $0.0997$
10% worst-case obj. $0\pm 0$ $0\pm 0$ $0\pm 0$
avg-case obj. $-1.98\pm 0.08$ $-1.51\pm 0.08$ $-1.83\pm 0.04$
@@
In Table 1, we compare the three uncertainty sets, with an average taken over the 10 independent runs. We note that while the 10% worst-case values are all 0, the average-case objective for the learned decision-focused uncertainty set (LRO) outperforms the other two methods, signifying that we attain less conservative - but still robust - solutions. In fact, we note that the worst average-case objective for LRO in the given (0.25,0.75) quantile range still outperforms the best average-case objective for the other methods. In the next sections, we introduce an automatic technique to find these contextual decision-focused sets. In the larger numerical examples in Section 6, we further demonstrate the ability of the method to outperform baseline methods in both worst-case and average-case performance.